Trico Company set the following standard unit costs for its single product. Direct materials (30 Ibs. @ $4.80 per Ib.) DirectActual costs incurred during the current quarter follow. Direct materials (1,582,000 lbs. @ $5.90 per $ 9,333,800 lb.) DirectRequired: 1. Compute the direct materials cost variance, including its price and quantity variances. Actual Cost Standard Cos2. Compute the direct labor cost variance, including its rate and efficiency variances. Actual Cost Standard Cost 3. ComputeFixed overhead volume variance Budgeted forced overhead Fixed overhead cost applied Ficed overhead volume variance
Answer #1

Solution 1:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
1582000 $5.90 $9,333,800.00 1582000 $4.80 $7,593,600.00 1593000 $4.80 $7,646,400.00
$1,740,200.00 U $52,800.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $1,740,200.00 U
Direct material quantity variance $52,800.00 F
Direct material cost variance $1,687,400.00 U

Solution 2:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
420800 $12.50 $5,260,000.00 420800 $16.00 $6,732,800.00 424800 $16.00 $6,796,800.00
$1,472,800.00 F $64,000.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $1,472,800.00 F
Direct Labor Efficiency variance $64,000.00 F
Direct labor cost variance $1,536,800.00 F

Solution 3:

Controllable Variance
Actual overhead $8,320,800.00
Budgeted overhead $8,354,400.00
Controllable variance $33,600.00 F
Fixed overhead volume variance
Budgeted fixed overhead $4,531,200.00
Fixed overhead cost applied $5,097,600.00
Fixed overhead volume variance $566,400.00 F